OpenStudy (anonymous):

Derivative of x(t) = 3 - |t - 3| ? i don't know how to find the derivative of an absolute value... help plz

OpenStudy (anonymous):

Derivative of absolute value is simple.. |x| = x, x > 0 -x, x < 0 Differentiation gives: d/dx |x| = 1, x > 0 -1, x < 0

OpenStudy (anonymous):

Use the chain rule on the rest of it.

OpenStudy (anonymous):

The function I provided actually is sometimes written as sign(x) It can also be written as x/|x|

OpenStudy (anonymous):

So you could use d/dx |x| = x / |x|

OpenStudy (vincent-lyon.fr):

Simple way is to rewrite the function without the absolute value for t>3 and then for t<3.

OpenStudy (amistre64):

if memory serves:\[\frac d{dx}|f(x)|\to\frac{f(x)}{|f(x)|}f'(x)\] \[|c-x|'\to\frac{c-x}{|c-x|}~(-1)\]

OpenStudy (anonymous):

@wio Thanks!